I am given 2, 3-dimensional rectangles. The dimensions on A are l=4, w=10, h=18. The dimensions on b are l=8, w=20, h=36. I have to find the perimeters and explain the relationship between changes in dimensions and perimeter of the base of rectanular solids. Using ratios in the explanation.
I've done P=2(lw) so for Rectangle A the perimeter = 80 and rectangle B the perimeter = 320. I see that all of the dimensions of A are 1/2 the dimensions of B and that the perimeter of of B is 4 times that of A. But how do I explain all of this using ratios?

Answer by venugopalramana(3286)

SEE MY COMMENTS BELOW.
--------------------------------------------------------------------------------
I am given 2, 3-dimensional rectangles.
THEY ARE CALLED CUBOIDS OR RECTANGULAR PARALLELOPIPEDS
The dimensions on A are l=4, w=10, h=18. The dimensions on b are l=8, w=20, h=36. I have to find the perimeters and explain the relationship between changes in dimensions and perimeter of the base of rectanular solids. Using ratios in the explanation.
I've done P=2(lw)
NO PERIMETER IS 2(L+W)FOR ANY SIDE.
so for Rectangle A the perimeter = 80
NO...THIS =2(4+10)=28
and rectangle B the perimeter = 320.
NO..THIS =2(8+20)=56
I see that all of the dimensions of A are 1/2 the dimensions of B and that the perimeter of of B is 4 times that of A.
NO PERIMETER IS ALSO 1/2 TIMES.
But how do I explain all of this using ratios?
YOU CONCLUDE THAT PERIMETER CHANGES IN SAME RATIO OR IN PROPORTION TO CHANGE IN RATIO OF LENGTH OF SIDES.